# GDPR Elements, and Future Directions

### v3

@@ I Introduction
The advent of Distributed Ledger Technology (DLT), and in particular blockchain, has become one of the most prominent phenomena of our time. As a result, it has come to the forefront of research in many disciplines (e.g., cryptocurrency, smart contracts, identity management) aimed at including distributed data storage. Due to the success of Bitcoin and cryptocurrencies, the corresponding blockchain technology is now successfully practiced to distribute data in a secure-by-design database without the need for a central authority. Not only have the unprecedented security properties of blockchain revolutionized a wide range of financial services and digital payments, but its unique characteristics such as decentralization, immutability, auditability, fault tolerance, and availability have played a pivotal role in receiving further public attention so far.
On the one hand, it is not surprising that, due to its availability and decentralization properties, the blockchain solves the security problems arising from storing biometric features (e.g., storing biometric features of the same individual in different places belonging to independent applications). In particular, in terms of biometric verification, the so-called on-chain storage of biometric templates helps to prevent Denial of Service (DoS) attacks caused by a single point of failure. However, a major breakthrough for the combination of blockchain and biometrics would be the introduction of the Public Key Infrastructure (PKI) mechanism[7]. Owing to their inherent connection to an individual’s identity, biometrics lack PKI structure. To address this limitation, the advantages of using blockchain and more specifically a novel introduced concept called distributed identifiers (DIDs) are being investigated. Although current studies are almost focused only on the role of the blockchain in biometric verification, identity management, and e-voting, we discuss the unprecedented merits of combining blockchain and biometrics for other potential applications such as using blockchain as computing power for biometrics or introducing a foundation for PKI mechanism. In addition, we discuss the role of blockchain in providing an infrastructure for multimodal biometrics or using it as a computational mean in Proof of Work (PoW)-based consensus for biometric verification.
@@
  • An application-based summary and categorization of the literature from the perspective of using the blockchain in favor of biometrics.
@@
In this regard, Section II gives a concise introduction to blockchain technology, biometrics, and their combination. In Section III, we analyze the impact of blockchain features when combined with biometrics from a technical point of view. This section discusses the prospects and limits that lie ahead of this combination. In this section, we also provide a categorization of the technical literature on combining blockchain and biometrics to improve biometric systems. Section IV contains the legal analysis based on GDPR. Section V offers some recommendations and suggestions that should be kept in mind for future studies in this line of research. Finally, Section VI concludes this survey.
@@ II-A Blockchain Technology
In particular, the header contains the following: the owner identifier and the identifier of the previous block, a timestamp denoting the generation time of the current block, a nonce, and a merkle tree root hash which denotes the hash value of all transactions stored in the block. The body consists of transactions and a transaction counter. The maximum number of transactions that fit within the body depends on the block size and the size of the transactions within. A digital signature based on asymmetric cryptography is used to validate transactions [11].
The most prominent characteristics of blockchain that have made it highly popular for integration with other technologies, including biometrics, are Decentralization, Immutability, Auditability, Fault tolerant, and Availability [12, 13].
@@ II-C Blockchain for Biometrics
A transaction on the blockchain represents an interaction between the parties [16]. In the context of biometrics, recording the result of an authentication is considered a transaction. Common biometric operations, such as template creation, modification, deletion, and retrieval, are also considered transactions that are recorded on the blockchain[17]. The lifestyle of a biometric transaction on the blockchain is shown in Fig. 2.
@@ II-D Related Works
Several studies are related to the current survey:
In 2022, Ahmed et al. [20] provided an in-depth analysis of how blockchain technology is revolutionizing identity management systems (IdMs), particularly through the lens of Self-Sovereign Identity (SSI). Their study explored the SSI ecosystem, emphasizing the role of decentralized identifiers and verifiable credentials in empowering users with control over their personal data.
@@
III Combining Blockchain and Biometrics: Literature
Biometric systems can greatly benefit from the functionality and features offered by the blockchain. In this section, we argue for the strengths and weaknesses of key aspects that play a pivotal role in the integration of blockchain and biometrics.
@@ III-A1 Storage
All blockchain-based applications face an inherent challenge with the size of blocks [25]. In existing scenarios, all transactions that are processed through the network have to be stored on the blockchain. In the blockchain network, all nodes are responsible for storing and maintaining the entire transaction record. With the current size and growing storage space of two prominent public blockchains, Bitcoin and Ethereum, exceeding 200GB [26], scalability has become the main barrier to adopting the blockchain in all applications, including biometrics. Multiple solutions have already been proposed, regardless of the application, to address this challenge. The first solution is to remove old recorded transactions and keep the balance of non-empty addresses using a database called the account tree [27]. This approach obviates the need for storing all transactions by keeping the hash of previous transactions to check the validity of new transactions.
The second solution would be to re-design the blockchain for the next generation of Bitcoin (Bitcoin-NG) [28]. The driving force behind this remedy is to divide conventional blocks into two separate parts: the key block and the microblock. In this revision, miners would compete for the key block to become the next leader who is eligible for generating a microblock wherein they store new transactions. Nevertheless, this solution suffers from some limits. For example, in terms of biometric verification, it needs to be lightweight due to the high utilization rate of this application. In case of biometric verification, if the history of all transaction are stored on-chain, it is essential to take care of storage obstacles. It is important to mention that it is considered risky by legal reasons to directly store raw biometric data or their extracted embeddings on a public blockchain. Thus, when we refer to storing biometric data, it usually means storing protected biometric templates. Lastly, the third solution, being considered by the most current literature [29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39] to overcome scalability issues with storage limitations [17], also applicable to biometrics, is to record blockchain transactions (biometric data) off-chain (any storage located outside of public blockchain). This solution which does not guarantee the immutability of the biometric data is divided into two approaches: (i) recording the hash of the biometric data (hash of the protected template) inside the public ledger and storing the biometric data itself off-chain in external storage systems, or (ii) recording only these hashes in a binary tree data structure called Merkle tree [26]. The key difference between these approaches lies in how they securely store content. The latter approach ensures data integrity through a recursive process of self-construction. The latest advancement in off-chain storage methods, designed to maintain the decentralization of public blockchains, is the use of IPFS [40]. The InterPlanetary File System (IPFS) is a distributed file system meant to be used for connecting computing devices with the same data structure. Like blockchain, it has no single point of failure where nodes do not need to trust each other. Yet, regarding the combination of blockchain and biometrics, most of the existing works suggest recording the hash of biometric data on-chain and storing the data itself on IPFS. Though this kind of storage introduces some deficiencies as the hash of data limits some biometric applications. For example, concerning biometric matching, extracted features are transformed into templates using different protection methods [41]. Therefore, storing the hash of templates inside a block makes it infeasible for biometric comparison function. Off-chain storage gives this opportunity to store the biometric data itself in addition to the hash of it. Hence, regarding biometric comparison, it can be done using a smart contract on off-chain networks. Currently, there are four recording approaches with regard to this combination in the existing literature: (i) Store biometric data as it is without any pre-processing on-chain (directly on the blockchain), (ii) Store biometric data off-chain on a side-chain (outside of the blockchain) and the hash of it on-chain, (iii) Store biometric data in distributed file systems (e.g., IPFS) and the hash of it on-chain, (iv) Store only the hash of biometric data on a Merkle tree. The pros and cons of each approach are given in Table I.
@@
Approach Pros Cons
@@ III-A2 Latency
Writing new blocks on-chain is quite slow and completely infeasible for being adopted in real-time applications. For example, in terms of the most popular cryptocurrency, the required time to establish a new block in the Bitcoin network is 10 minutes on average, making it limited to processing up to 7 transactions per second. The same scalability limit lies ahead of the second generation of blockchain systems such as Ethereum with a limit of 15 transactions per second. As a result, biometric applications that require real-time responses, such as most verification and identity management systems, cannot integrate with these blockchain networks. Off-chain storage has also been proposed to solve the latency problem [42]. In this case, IPFS can address the content quicker, more reliably, and more easily by locating the stored data using the file information. Although off-chain storage helps reduce latency compared to on-chain solutions, it still imposes limitations on certain biometric applications (e.g., biometric authentication, surveillance, e-commerce), particularly similar to the storage challenges previously mentioned.
@@ III-A3 Computation cost
It is discussed that another reason to suggest using off-chain storage is attributable to the total cost of the computation on-chain. The main biometric operations that impose arithmetic costs on the blockchain network involve setup, template transformation, encryption, decryption, upload (read, write), updating the smart contract, and calculating distance. However, the exact computational cost of biometrics in the blockchain is highly dependent on the template acquisition method [32], the underlying consensus protocol, and the operations executed by the contract [30]. For example, the initial cost for reading and writing 1KB of data on the Ethereum network is 6400 and 640000 gas units, respectively [17]. Therefore, storing 1KB of data from the vector of face characteristics for the VGG-Face template on the Ethereum network costs 352912 gas. Gas is a unit to pay for transactions on the Ethereum blockchain denoted in gwei, where 1 ETH = $1*10^{9}$ (1,000,000,000) gwei. It is a measure to compute how much a transaction would cost. Therefore, at the time of writing this article, the storage cost for 1KB data of the face feature vector on a full on-chain storage scheme is approximately \$0.69 (1 ETH = \$1974). Due to the computationally intensive process of biometric applications such as biometric verification, the widely suggested solution for massive scalability appears to be to resort to off-chain storage (approaches ii and iii in Table I), regardless of the limitations.
@@ III-C Consensus
In the blockchain context, a consensus is a mechanism to reach an agreement between participants who do not necessarily trust each other in a distributed network without a central authority. The consensus protocol is responsible for the correct execution of the smart contract on the blockchain. In combination with biometrics, the consensus algorithm affects the latency to execute operations. With the emergence of the next generation of blockchain networks [42], equipped with fast consensus algorithms that maintain the security benefits of their predecessors, the need for off-chain protocols, at least to overcome scalability limits, will likely be eliminated. To this end, the third generation of blockchain systems is emerging that targets scalability limitations. As an example of these new generations of consensus algorithms, Stellar consensus protocol [45], which is based on a new type of federated byzantine agreement [46], can process 1,000 transactions/second, Ripple can handle up to 1,500 transactions/second, and Hashgraph, which takes advantage of a new type of asynchronous, byzantine fault tolerance can achieve above 250,000 transactions/second. Recently, an identity management method called DID (Distributed Identifiers) [47] has been introduced which transfers computation overload off-chain, decreasing demands for the main blockchain. Due to its decentralized nature and ability to enable trusted interactions, DID serves as a viable alternative for storing biometric templates on-chain. We will discuss DID further in Section V-B.
@@ III-D Security Enhancement
Owing to its unique features, blockchain can potentially help improve the security of distributed networks. In terms of the combination with biometrics, while it offers a breakthrough, multiple concerns need to be addressed. First and foremost, this integration can not only help introduce a public key infrastructure (PKI) platform [7] to biometrics, but also helps to overcome the problem of a single point of failure with which conventional PKIs are entangled. In conventional PKI systems, if the Certificate Authority (CA) is not accessible, it is not possible to verify the authenticity of generated certificates. Due to its availability characteristics, blockchain addresses this shortcoming. Second, considering biometric verification, it provides a distributed storage system in which multiple transformations of the same biometric template can be stored in different places [48, 49]. In existing scenarios, some tried to increase blockchain security by generating biometric cryptographic keys to secure the private key in blockchain wallets [50] or creating an encryption scheme to sign blockchain transactions [51], while others took advantage of the blockchain to increase biometric system security. Regarding the first, since the private key is stored in the user’s wallet, the security of the private key is dependent on the security of the wallet password. Biometrics can also be used as an access control to the wallet in a key binding approach [52]. In terms of the latter, which is the focus of this survey, most studies have been carried out to secure biometric verification using blockchain as a means to provide decentralized auditability, availability, and immutability. Some studies suggest blockchain for biometric applications where transparency is required, such as public data management, electronic voting, and taxation [39, 35]. Since the blockchain is immutable and the recorded information persists there forever, the fate of stored biometric templates would be shifted to the security of the underlying cryptographic method. Concerning auditability, referring to the 51% attack [53], it is possible to exert influence by most of the nodes that colluded for a single purpose and contradict the facts already recorded by creating a fork. Furthermore, due to the immutability of the blockchain, it is not possible to discard an expired biometric template stored earlier on a public ledger unless $51\%$ of the nodes colluded. Regarding private blockchains, since the owner of the blockchain controls the majority ($51\%$) of the miners, it is practically possible to remove an expired biometric template by resetting the last block to an earlier block and continuing to mine from there. However, this process is costly, since removing a biometric template from the blockchain also deletes all subsequently added templates. Thus, the owner of the private blockchain needs to rewrite the history after the discarded block. A recently proposed solution for Bitcoin introduces a miner-redactable and scalable approach that reportedly eliminates the need for trustees, heuristics, or hard forks. Kiraz et al. [54] claim to achieve instant erasure of all transaction types by enforcing transaction “templates”. These templates ensure that only the non-executable portions of unlocking scripts can be modified, preventing unintended alterations. In particular, a significant portion of such nonexecutable data consists of nonfinancial information such as biometric templates or references to them within the Bitcoin network, which are stored as OP_RETURN data. In 2017, reports indicate that 86.8% of Bitcoin’s non-financial data falls into this category. Furthermore, this approach eliminates the need for modifications in child transactions after redaction, effectively resolving the issue of “cascade of changes” found in $\mu$Chain [55], a redactable blockchain designed for Ethereum.

### v4

# GDPR Analysis, and Future Directions

@@ I Introduction
The advent of Distributed Ledger Technology (DLT), and in particular blockchain, has become one of the most prominent phenomena of our time. As a result, it has come to the forefront of research in many disciplines (e.g., cryptocurrency, smart contracts, identity management) aimed at including distributed data storage. Following the success of Bitcoin and other cryptocurrencies, the underlying blockchain technology is now widely used to distribute data through secure-by-design databases without the need for a central authority. Not only have the unprecedented security properties of blockchain revolutionized a wide range of financial services and digital payments, but its unique characteristics such as decentralization, immutability, auditability, fault tolerance, and availability have played a pivotal role in receiving further public attention so far.
On the one hand, it is not surprising that, due to its availability and decentralization properties, the blockchain solves the security problems arising from storing biometric features (e.g., storing biometric features of the same individual in different places belonging to independent applications). In particular, in terms of biometric verification, the so-called on-chain storage of biometric templates helps to prevent Denial of Service (DoS) attacks caused by a single point of failure. However, a major breakthrough for the combination of blockchain and biometrics would be the introduction of the Public Key Infrastructure (PKI) mechanism[7]. Owing to their inherent connection to an individual’s identity, biometrics lack PKI structure. To address this limitation, the advantages of using blockchain and more specifically a novel introduced concept called distributed identifiers (DIDs) are being investigated. Although current studies are almost focused only on the role of the blockchain in biometric verification, identity management, and e-voting, we discuss the unprecedented merits of combining blockchain and biometrics for other potential applications such as using blockchain as computing power for biometrics or introducing a foundation for PKI mechanism. In addition, we discuss the role of blockchain in enabling multimodal biometrics, providing self-Sovereign Identity (SSI), and serving as a computational medium in Proof-of-Work (PoW)-based consensus for biometric verification.
@@
  • A summary and categorization of the literature from the perspective of key factors in applying blockchain to biometrics.
@@
In this regard, Section II gives a concise introduction to blockchain technology, biometrics, and their combination. In Section III, we analyze the impact of blockchain features when combined with biometrics from a technical point of view. This section discusses the prospects and limits that lie ahead of this combination. In this section, we also provide a categorization of the technical literature on combining blockchain and biometrics to improve biometric systems. Section IV contains the legal analysis based on GDPR. Section V offers some recommendations and potential research directions for future studies in this field. Finally, Section VI concludes this survey.
@@ II-A Blockchain Technology
In particular, the header contains the following: the owner identifier and the identifier of the previous block, a timestamp denoting the generation time of the current block, a nonce (a number used in PoW to find a valid block hash, or in transactions to ensure uniqueness and ordering), and a merkle tree root hash which denotes the hash value of all transactions stored in the block. The body consists of transactions and a transaction counter. The maximum number of transactions that fit within the body depends on the block size and the size of the transactions within. A digital signature based on asymmetric cryptography is used to validate transactions [11].
The most prominent characteristics of blockchain that have made it highly popular for integration with other technologies, including biometrics, are decentralization, immutability, auditability, fault tolerance, and availability [12, 13].
@@ II-C Blockchain for Biometrics
A transaction on the blockchain represents an interaction between the parties [16]. In the context of biometrics, recording the result of an authentication is considered a transaction. Common biometric operations, such as template creation, modification, deletion, and retrieval, are also considered transactions that are recorded on the blockchain[17]. The lifecycle of a biometric transaction on the blockchain is shown in Fig. 2.
@@ II-D Related Works
Several studies are related to the current survey. In 2022, Ahmed et al. [20] provided an in-depth analysis of how blockchain technology is revolutionizing identity management systems (IdMs), particularly through the lens of SSI. Their study explored the SSI ecosystem, emphasizing the role of decentralized identifiers and verifiable credentials in empowering users with control over their personal data.
@@
III Blockchain and Biometrics: Advantages and Limitations
Biometric systems can greatly benefit from the functionality and features offered by the blockchain. Fig. 4 presents our proposed taxonomy of blockchain features applied to biometric integrations, based on existing works. In this section, we discuss the strengths and weaknesses of key aspects that play a pivotal role in the integration of blockchain and biometrics.
Figure 4: The Proposed Taxonomy of Current Blockchain–Biometric Integration Approaches
@@ III-A1 Storage
All blockchain-based applications face an inherent challenge related to the limited block size and the continuous growth of stored data [25]. In such systems, every transaction that is processed through the network must be permanently recorded on the blockchain, and all participating nodes are responsible for maintaining the complete transaction history. With the current size and growing storage space of two prominent public blockchains, Bitcoin and Ethereum, exceeding 200GB [26], scalability has become the main barrier to adopting the blockchain in all applications, including biometrics. Broadly, three categories of solutions have been proposed:
The first solution is to remove old recorded transactions and keep the balance of non-empty addresses using a database called the account tree [27]. This approach obviates the need for storing all transactions by keeping the hash of previous transactions to check the validity of new transactions.
The second solution involves redesigning blockchain protocols, such as Bitcoin-NG [28]. This approach divides conventional blocks into two parts: a key block and a microblock. Miners compete for the key block to become the next leader, who then generates microblocks to store new transactions. However, this design has limitations. For biometric verification, it must remain lightweight due to high usage rates, and storing all transactions on-chain raises storage concerns. Moreover, storing raw biometric data or embeddings directly on a public blockchain is legally risky, so typically only protected biometric templates are stored.
Lastly, the third solution, widely discussed in the recent literature [29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40] is off-chain storage, which records biometric transactions outside the blockchain while maintaining cryptographic references on-chain. This approach, though it does not guarantee full immutability, offers the most practical scalability benefits. It is typically implemented either by (i) recording the hash of biometric data within the ledger while storing the data itself in an external repository, or (ii) recording only these hashes within a Merkle tree [26], which ensures integrity through a recursive process of self-construction.
A more advanced form of off-chain storage that preserves decentralization is the InterPlanetary File System (IPFS) [41], a distributed file system connecting devices via a shared data structure. Like the blockchain itself, IPFS has no central authority or single point of failure. Most blockchain–biometric studies propose storing the hash of biometric data on-chain and the data itself on IPFS. However, this setup limits direct biometric matching, since hashes cannot be used for comparison. Biometric verification thus requires the protected template to remain accessible off-chain, where matching can be performed through smart contracts or auxiliary verification modules.
In summary, four main recording strategies exist for combining blockchain and biometrics: (i) storing biometric data directly on-chain, (ii) storing data off-chain on a side-chain with its hash on-chain, (iii) storing data in distributed file systems (e.g., IPFS) with on-chain hashes, and (iv) storing only hashes on a Merkle tree. The advantages and disadvantages of each approach are summarized in Table I.
@@
Approach Advantages Limitations
@@ III-A2 Latency
Writing new blocks on-chain is quite slow and completely infeasible for being adopted in real-time applications. For example, in terms of the most popular cryptocurrency, the required time to establish a new block in the Bitcoin network is 10 minutes on average, making it limited to processing up to 7 transactions per second. The same scalability limit lies ahead of the second generation of blockchain systems such as Ethereum with a limit of 15 transactions per second. As a result, biometric applications that require real-time responses, such as most verification and identity management systems, cannot integrate with these blockchain networks. To address this issue, next-generation blockchain networks overcoming these limitations should be developed. Off-chain storage has also been proposed to mitigate the latency problem [42]. In this case, IPFS can retrieve the stored content more quickly and reliably by locating it through its file information. Although off-chain storage helps reduce latency compared to on-chain solutions, it still imposes limitations on certain biometric applications (e.g., biometric authentication, surveillance, e-commerce), particularly similar to the storage challenges previously mentioned.
@@ III-A3 Computation cost
It is discussed that another reason to suggest using off-chain storage is attributable to the total cost of the computation on-chain. The main biometric operations that impose arithmetic costs on the blockchain network involve setup, template transformation, encryption, decryption, upload (read, write), updating the smart contract, and calculating distance. However, the exact computational cost of biometrics in the blockchain is highly dependent on the template acquisition method [32], the underlying consensus protocol, and the operations executed by the contract [30]. For example, the initial cost for reading and writing 1KB of data on the Ethereum network is 6,400 and 640,000 gas units, respectively [17]. Therefore, storing 1 KB of data from the vector of face characteristics, for instance from the VGG-Face template, on the Ethereum network costs approximately 352,912 gas. Gas is a unit to pay for transactions on the Ethereum blockchain denoted in gwei, where 1 ETH = $1*10^{9}$ (1,000,000,000) gwei. It is a measure to compute how much a transaction would cost. Therefore, at the time of writing this article, with a gas price of $0.635$ gwei per gas, the cost of storing 1 KB of face feature vector data on a full on-chain storage scheme is approximately \$1.03 (1 ETH = \$4,596.57). Due to the computationally intensive process of biometric applications such as biometric verification, the widely suggested solution for massive scalability appears to be to resort to off-chain storage (approaches ii and iii in Table I), regardless of the limitations.
@@ III-C Consensus
In the blockchain context, consensus is a mechanism to reach an agreement between participants who do not necessarily trust each other in a distributed network without a central authority. The consensus protocol is responsible for the correct execution of the smart contract on the blockchain. In combination with biometrics, the consensus algorithm affects the latency to execute operations. With the emergence of the next generation of blockchain networks [42], equipped with fast consensus algorithms that maintain the security benefits of their predecessors, the need for off-chain protocols, at least to overcome scalability limits, will likely be eliminated. To this end, the third generation of blockchain systems is emerging that targets scalability limitations. As an example of these new generations of consensus algorithms, Stellar consensus protocol [45], which is based on a new type of federated byzantine agreement [46], can process 1,000 transactions/second, Ripple can handle up to 1,500 transactions/second, and Hashgraph, which takes advantage of a new type of asynchronous, byzantine fault tolerance can achieve above 250,000 transactions/second. Recently, an identity management method called DID (Distributed Identifiers) [47] has been introduced which transfers computation overload off-chain, decreasing demands for the main blockchain. Due to its decentralized nature and ability to enable trusted interactions, DID serves as a viable alternative for storing biometric templates on-chain. We will discuss DID further in Section V-B.
@@ III-D Security Enhancement
Owing to its unique features, blockchain can potentially help improve the security of distributed networks. In terms of the combination with biometrics, while it offers a breakthrough, multiple concerns need to be addressed.
First and foremost, this integration can not only help introduce a public key infrastructure (PKI) platform [7] to biometrics, but also helps to overcome the problem of a single point of failure with which conventional PKIs are entangled. In conventional PKI systems, if the Certificate Authority (CA) is not accessible, it is not possible to verify the authenticity of generated certificates. Due to its availability characteristics, blockchain addresses this shortcoming. Second, considering biometric verification, it provides a distributed storage system in which multiple transformations of the same biometric template can be stored in different places [48, 49]. In existing scenarios, some tried to increase blockchain security by generating biometric cryptographic keys to secure the private key in blockchain wallets [50] or creating an encryption scheme to sign blockchain transactions [51], while others took advantage of the blockchain to increase biometric system security.
Regarding the former, because the private key resides in the user’s wallet, its security depends on the strength of the wallet’s password. Biometrics can also be used as an access control to the wallet in a key binding approach [52]. As for the latter, our focus in this survey, most studies try to secure biometric verification by leveraging blockchain for decentralized auditability, availability, and immutability. Some studies suggest blockchain for biometric applications where transparency is required, such as public data management, electronic voting, and taxation [39, 35].
Since the blockchain is immutable and the recorded information persists there forever, the fate of stored biometric templates would be shifted to the security of the underlying cryptographic method. Concerning auditability, referring to the 51% attack [53], it is possible to exert influence by most of the nodes that colluded for a single purpose and contradict the facts already recorded by creating a fork. Furthermore, due to the immutability of the blockchain, it is not possible to discard an expired biometric template stored earlier on a public ledger unless $51\%$ of the nodes colluded. Regarding private blockchains where a single organization controls the validator set, it is practically possible to remove an expired biometric template by reverting the ledger to a prior state (or producing an agreed fork) and then producing new blocks from that point. However, this process is costly, since removing a biometric template from the blockchain also deletes all subsequently added templates. Thus, the owner of the private blockchain needs to rewrite the history after the discarded block.
[…]